Constant-boiling HCl can be used as a primary standard for acid-base titrations. A 50.0 mL sample of constant-boiling HCl with a concentration of 0.1166 M was collected and titrated to an end point with 37.17 mL of Ba(OH)2 solution.
 
What is the molarity of the Ba(OH)2 solution?
Answer #1
Concepts and reason

Molarity (M) is calculated as the number of moles (n) divided by the volume (V).

M=nVM = \frac{n}{V}

1 mole consists of 6.022×10236.022 \times {10^{23}} atoms. This is called Avogadro’s Number.

Fundamentals

The reaction between hydrochloric acid and Barium hydroxide is as follows:

2HCl+Ba(OH)2BaCl2+2H2O{\rm{2HCl + Ba}}{\left( {{\rm{OH}}} \right)_{\rm{2}}}\longrightarrow{{}}{\rm{BaC}}{{\rm{l}}_{\rm{2}}}{\rm{ + 2}}{{\rm{H}}_{\rm{2}}}{\rm{O}}

Calculate the number of moles of HCl (nHCl)\left( {{n_{{\rm{HCl}}}}} \right) from the volume of HCl (V) and concentration of HCl (c)as follows:

nHCl=Vc{n_{{\rm{HCl}}}} = Vc

Substitute V as 50 mL and c as 0.1166 M in the above expression as follows:

nHCl=(50mL)(1L1000mL)(0.1166M)=0.00583mol\begin{array}{c}\\{n_{{\rm{HCl}}}} = \left( {50{\rm{ mL}}} \right)\left( {\frac{{1{\rm{ L}}}}{{1000{\rm{ mL}}}}} \right)\left( {0.1166{\rm{ M}}} \right)\\\\ = 0.00583{\rm{ mol}}\\\end{array}

Calculate the number of moles of Barium hydroxide (nBa(OH)2)\left( {{n_{{\rm{Ba}}{{\left( {{\rm{OH}}} \right)}_2}}}} \right) as follows:

nBa(OH)2=12nHCl{n_{{\rm{Ba}}{{\left( {{\rm{OH}}} \right)}_{\rm{2}}}}} = \frac{1}{2}{n_{{\rm{HCl}}}}

Substitute nHCl{n_{{\rm{HCl}}}} as 0.00583 mol in the above expression as follows:

nBa(OH)2=12(0.00583)mol=0.002915mol\begin{array}{c}\\{n_{{\rm{Ba}}{{\left( {{\rm{OH}}} \right)}_{\rm{2}}}}} = \frac{1}{2}\left( {0.00583} \right){\rm{ mol}}\\\\ = 0.002915{\rm{ mol}}\\\end{array}

Molarity of Barium hydroxide is calculated as follows:

M=nBa(OH)2VM = \frac{{{n_{{\rm{Ba}}{{\left( {{\rm{OH}}} \right)}_2}}}}}{V}

Substitute nBa(OH)2{n_{{\rm{Ba}}{{\left( {{\rm{OH}}} \right)}_2}}} as 0.002915 mol and V as 37.17 mL in the above expression as follows:

M=0.002915mol37.17mL(1L1000mL)=0.002915mol0.03717L=0.07842M\begin{array}{c}\\M = \frac{{0.002915{\rm{ mol}}}}{{37.17{\rm{ mL}}\left( {\frac{{1{\rm{ L}}}}{{{\rm{1000 mL}}}}} \right)}}\\\\ = \frac{{0.002915{\rm{ mol}}}}{{0.03717{\rm{ L}}}}\\\\ = 0.07842{\rm{ M}}\\\end{array}

Ans:

The molarity of Ba(OH)2{\rm{Ba}}{\left( {{\rm{OH}}} \right)_{\rm{2}}} solution is 0.07842 M.